To study of effect of physician participation in administrative decision making in hospitals, a random sample of 200 physicians were drawn from different hospitals in Arkansas. Each physician was asked whether he/she generally approved of the administrative decisions made within the hospital and whether they had active participation in the decision process. The results of the interviews are shown in table below. To examine the association between physician’s approval status of decision and their participation status in decision making, a chi-square test was used. Find the degrees of freedom.

Participate Not Participate
Approve 73 51
Do no approve 27 49
Answer #1

Solution :

From given that ,

r = 2

c = 2

degrees of freedom = (r - 1) * (c - 1) = (2 - 1) * (2 - 1) = 1